A gas occupies a volume of 845.4 mL of at standard temperature. If the temperature increases to 20.1oC, what is the new volume?


1
Expert's answer
2021-02-10T01:32:34-0500

According to the ideal gas law of the isobaric process:

V1/T1 = V2/T2

where V1 - initial volume, V2 - final volume, T1 - initial temperature, T2 - final temperature (20 °C = 293.25 K).

V2 = V1T2 / T1 = 845.4 mL × 293.25 K / 273.15 K = 907.6 mL


Answer: 907.6 mL